Ca(CO3)(aq) + 2HCl(aq) → CaCl2(aq) + H2O(l) + CO2(g)

If 0.0600 g HCl reacts with excess Ca(CO3), what amount (in moles) of CaCl2 will be produced?
1
Expert's answer
2020-10-19T14:26:53-0400

Solution:

The balanced chemical equation:

CaCO3(aq) + 2HCl(aq) → CaCl2(aq) + H2O(l) + CO2(g)

According to the equation: n(HCl) / 2 = n(CaCl2)

Moles of CaCl2 = Moles of HCl / 2


Moles of HCl = Mass of HCl / Molar mass of HCl

The molar mass of HCl is 36.46 g/mol.

Moles of HCl = (0.0600 g) / (36.46 g/mol) = 0.001646 mol


Hence,

Moles of CaCl2 = (0.001646 mol) / 2 = 0.000823 mol


Answer: 0.000823 moles of CaCl2 will be produced.